vidyo

New vidyo room:

Action Items:

News:

Franco: at the CB last week the new SP has chosen the team (deputies Malgeri, McBride), (physCoord W. Adam) Discussion ongoing on how to elect the SP. Right now it is 1 vote per institution. A committee has been formed. Long discussion on the upgrade coordinator. No decision whether to keep Didier or not in the role.

Framework & production

Analysis

B physics

Stefano: accepted by PLB, I swear I will bring prosecco. Reviewer happy about the details provided, understood the case is different from LHCb, accepted the rational to fit w/o the s-wave. The uncertainty in the s-wave is compatible with the expected contribution (~10%), the fit converged and unc on the results are compatible.

Charged Higgs to top-beauty

Battling with the smoothing. Trying w/ different options. Probably will need to disentangle all the contributions on the JES unc. Presenting results this or next week.

Dibosons, VZ resonance (V=W,Z)

There was the effort to combine 2016 results. Zucchetta is trying to coordinate the effort. In about 1-2 months the results should be ready.

4B full hadronic

Andres: got the Q from the ARC. In a couple of days we should have the green light and move to approval. This with the single SM training. The discrepancy is at ~2sigma.

Heavy New Physics

Matteo will write here!

VH SM

Embrional...

Same sign dileptons

Ugo: repeated the analysis on legacy NanoAOD, followed the prescriptions. The excess is reduced because the 3 methods before were having an agreement which was better. Now the agreement is worse btw two of them, so the syst are bigger. When the full datasample will be available the third method should be the best one from the systematics point of view. Now we start seeing effects also in other distributions, and 2D plots. This is visible in 2016-2017 data. Need to have the MC for the 2017 still.

EPR

Triggerless

Chat with the DAQ ppl at CERN. Meyer was there too. Presented the status, well received. Worked with Andrea. The 40MHz readout were moved to legnaro. Friday performed the installation. Meyer suggests to try to see whether we can the L1T can be also intercepted and process them at 40MHz, for monitoring, and maybe analysis. A place would be the input to the global mu trigger, where the info from the CALO is arriving. This would be a Run3 test.

Detector

DT

Phase 1 validation work for the student (Luca Bonaldo). The idea is to coordinate with Bachtis and Costas to understand whether we can contribute. It seem that the work of Nicola can fit well in the line of development of Bachtis. GIF++ test beam starting tomorrow.

Luminosity & Bril

Paul: improvements in systematics. Uncertainty reduced. We had stable beam. Tested the linearity of luminosity. Luminometers linear, discrepancy among luminometers are still there.

Timing Layer

Tracker

A.O.B.:

-- Main.Franco Simonetto 04-10-2017 -- FrancoSimonetto - 2018-03-16

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2018-04-23 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback